OpenStudy (anonymous):

Complete the square to write in vertex form: y=3x^2-6x+5

OpenStudy (anonymous):

3(x^2-3x)+5=3(x^2-3x+9/4)-27/4=3(x-3/2)^2-27/4 Then the vertex is (3/2,-27/4)